OpenStudy (anonymous):

I am on the verge of a panic attack, i missed 3 weeks of school because i was sick. I am still sick but im not throwing up anymore so i am attending school. I hae finals tomorrow and i need everything to be able to get ready! I have a lot of questions and not much time, so it would be the quickest way if people could answer and explain. I am going to be up all night and i need help!!!!!!

OpenStudy (anonymous):

Post some stuff!

OpenStudy (anonymous):

for \[f(x) = \frac{ x ^{2} }{ x-1 }\] use the definition of differentiability \[f \prime (c) \frac{ f( x +\Delta x) - f(x) }{ \Delta x }\] to determine if f(x) is differentiable at x=1

OpenStudy (anonymous):

@stardogchampion

OpenStudy (anonymous):

hi wat grade are u in

OpenStudy (anonymous):

12th

OpenStudy (anonymous):

wat school

OpenStudy (anonymous):

im taking finals 2

OpenStudy (anonymous):

hello

OpenStudy (anonymous):

any1 here

OpenStudy (anonymous):

it can't be differentiable at 1 because the function is not even defined at 1

OpenStudy (anonymous):

this is clear right? if the function does not exist somewhere, obviously it's derivative doesn't

OpenStudy (anonymous):

i think she left

OpenStudy (anonymous):

yes it is thank you! do you have time to help me more???

OpenStudy (anonymous):

sure

OpenStudy (anonymous):

btw we did not bother to actually compute \(f \prime (c)= \frac{ f( x +\Delta x) - f(x) }{ \Delta x }\) because it is a stupid idea, but we can try it if you like and see why it does not exist

OpenStudy (anonymous):

i don't see the point in that if we are going to get the same answer. can you explain why functions with corners are not differentiable even though it is continuous?

OpenStudy (anonymous):

which in any case should be \[f \prime (x) =\frac{ f( x +\Delta x) - f(x) }{ \Delta x }\]or \[f \prime (c) =\frac{ f( c +\Delta x) - f(c) }{ \Delta x }\]

OpenStudy (anonymous):

yes i think so

OpenStudy (anonymous):

intuitively it is because the derivative evaluated at a number is the slope of the line tangent to the graph at that number if the function has a corner at that number, there is no definition for one tangent line as in this picture 

OpenStudy (anonymous):

reason two the derivative is a limit in order for a limit to exist it has to exist from the left, and from the right, and those limits have to be equal in the famous example of why \(f(x)=|x|\) does not have a derivative at the corner \((0,0)\) is because the limit of the difference quotient from the left is \(-1\) and from the right is \(1\) and those number are not the same 

OpenStudy (anonymous):

okay thank you.

OpenStudy (anonymous):

yw

OpenStudy (anonymous):

Use the following graph of f(x) to describe the behavior of f ‘(x). State where f ‘(x) is positive, negative, zero or not defined.

OpenStudy (anonymous):

is it positive? or not defined?

OpenStudy (anonymous):

it is not defined where the function is not defined, at \(x=2\)

OpenStudy (anonymous):

so any with an asymptote is not defined?

OpenStudy (anonymous):

the derivative is the slope of the tangent lines the slopes will be positive over the intervals where the function is increasing, i.e. going up

OpenStudy (anonymous):

function doesn't exist at the vertical asymptote, neither does the derivative

OpenStudy (anonymous):

derivative will be negative over the interval where the function is decreasing

OpenStudy (anonymous):

okay, i get that

OpenStudy (anonymous):

looks like your function is increasing over \((-3,2)\cup (3,5)\) and decreasing on \((2,3)\)

OpenStudy (anonymous):

derivative will be zero where the tangent line is horizontal, if i am reading the graph correctly looks like it will be horizontal as \(x=3\)

OpenStudy (anonymous):

a good candidate for a local maximum or a local minimum in this case \(x=3 \) looks to be a local min

OpenStudy (anonymous):

Use the following graph of f (x) to find the graph of f ‘(x).

OpenStudy (anonymous):

that function is decreasing, then increasing look for a derivative that is negative (below the x axis) then positive (above the x axis)

OpenStudy (anonymous):

i wanna say it is the last one?

OpenStudy (anonymous):

that is a tough one i might say 143 if you look at the rate of decrease of the original function, over one interval it looks like it is decreasing at a constant rate if you look at the derivative 143, you see it is almost horizontal over the same interval

OpenStudy (anonymous):

gotta run, back later post again if you have more questions in a new thread

OpenStudy (anonymous):

kk thank you!